A surface encloses the charges q1 = 3.1 µC, q2 = 6.9 µC, and q3 = -3.2 µC. Find the electric flux through this surface.
N·m2/C
Answer #1

Electric flux = E . da = q/ _0

                                = (3.1 + 6.9 - 3.2) * 10-6 / 8.85 * 10-12 = 0.768 * 106   N m2 / c